A child with a mass of 30 kg is standing on a spring scale in an elevator. If the spring scale reads 360 N, what are the magnitude and direction of the acceleration of the elevator at this time? The acceleration due to gravity is 10 m/s2.
Answer #1


apply F = m(a+g)

ma = 360-30 *10

ma = 60

a = 60/30

a = 2 m/s^2
Answer #2

real weight =mg=30*10=300N
given weight =360N

given weight>real weight

so the elevator must be moving in downward direction

m(g+a)=360

30(10+a)=360

acceleration =2m/sec^2 ,downward direction